OpenStudy (asib1214):

Determine the number of moles of ammonia that are produced when 43.5g of nitrogen reacts with an excess of hydrogen. Balance the equation for this reaction. Thanks in advance!!!!

OpenStudy (abmon98):

N2+3H2-->2NH3 Number of moles=mass/molar mass N=43.5/28 N=1.55 moles Ratio of moles of nitrogen to ammonia is 1:2 so if 1.55 moles of nitrogen were consumed we 1.55*2 are produced of ammonia gas

OpenStudy (asib1214):

so it'll be N2 + 3H2 = 2NH3 43.5g 28.2g/mol number of moles= 43.5X 1mol/28.2g=1.54moles.

OpenStudy (abmon98):

thats the number of moles of nitrogen

OpenStudy (asib1214):

so number of moles of ammonia would be like this.... 1.54molN2 X 2molNH3/1molN2= 3.01moles right

OpenStudy (abmon98):

exactly

OpenStudy (asib1214):

aight thanks bro, we need people like you here on open study. GOD bless you!!!! cheers

OpenStudy (asib1214):

just one quick question, according to the calculations, we get 3 moles of NH3 but when you balance the equation we get 2 moles? why is that

OpenStudy (abmon98):

your most welcome :D i really appreciate that :. okay let me try 3 moles of ammonia if we had 3 moles of ammonia then we would have 9/2 moles of hydrogen consumed and 1.5 moles of nitrogen if we used the same method of calculation would give the exact value when we used 2 moles of ammonia instead. you may workout with fraction if thats easy for you :)

OpenStudy (asib1214):

thanks and sorry for being a flatout dumbass for a second. i've figured out what's going on. Again thanks a mole.

OpenStudy (abmon98):

sorry for what you are free to ask on whatever your stuck, stop saying that your being a dumbass believe in yourself and good to hear that you have figured out how to solve such problem hope its clear now :D